Sally buys a radio for $120. She pays for it with $2 and $5 bills. If Sally used a total of 30 bills, how many of each did she use?
The unit is in solving problems using one variable. What would the algebraic equation be for this problem?

Answer by stanbon(48558)

Sally buys a radio for $120. She pays for it with $2 and $5 bills. If Sally used a total of 30 bills, how many of each did she use?
The unit is in solving problems using one variable. What would the algebraic equation be for this problem?
--------------------
Let # of $5 bills be "x"; Value of these is 5x dollars
----------------------
# of $2 bills is "30-x"; value of these is 2(30-x)= 60-2x dollars
----------------------
EQUATION:
value + value = $120
5x + 60-2x = 120
3x = 60
x = 20 (# of $5 bills)
30-x = 10 (# of $2 bills)
=================
Cheers,
Stan H.